OpenStudy (anonymous):

Solve V = π r^2h for h.

OpenStudy (anonymous):

Divide both sides by πr².

OpenStudy (anonymous):

Vh=r^2h (Vh)/r^2 =h so, h=\[\frac{ Vh }{{{r}^2} }\]

OpenStudy (anonymous):

http://openstudy.com/study#/updates/507acfa2e4b035d93135a7f4 @Bhagyashree